OpenStudy (anonymous):

I have an indirect interval question. The problem tells me to evaluate the intergral x^3(x^4+5)^-1/2. getting stuck on the alegbra.............

myininaya (myininaya):

let u=x^4+5 so du=4x^3 dx or du/4=x^3dx

myininaya (myininaya):

so now you have 1/4 * int(u^(-1/2))

OpenStudy (anonymous):

Yikes, ok. I wasn't aware I had to use that rule. thank you